# Emilia: An Extensive, Multilingual, and Diverse Speech Dataset for Large-Scale Speech Generation

### v2

@@ 4.3 Multilingual TTS
En SoundStorm 6.2% 0.614 14.82
VoiceBox 5.8% 0.581 14.47
Zh SoundStorm 4.1% 0.564 35.59
VoiceBox 4.7% 0.560 48.95
De SoundStorm 6.8% 0.680 32.72
VoiceBox 13.3% 0.633 44.68
Fr SoundStorm 8.2% 0.589 42.92
VoiceBox 17.5% 0.550 48.82
Ja SoundStorm 3.6% 0.625 49.42
VoiceBox 10.8% 0.562 49.39
Ko SoundStorm 10.9% 0.681 47.93
VoiceBox 15.2% 0.608 63.00
Next, we conduct experiments to evaluate the performance of the multilingual TTS capability of the baselines trained on the full Emilia datasets, encompassing six languages. The test set for En is Emilia-Test. The test set for Zh is from Aishell-3. The test sets for De, Fr, Ja, and Ko are from Common Voice. Each test set contains at least 500 samples. Due to space limitations, we only present the objective evaluation results in Table 5, omitting the subjective evaluation results. The findings confirm that both models exhibit strong zero-shot multilingual TTS performance, underscoring the multilingual effectiveness of the Emilia dataset.
@@ 5 Conclusions
In conclusion, this paper introduces Emilia, an extensive, multilingual, and diverse dataset for speech generation, along with Emilia-Pipe, an open-source preprocessing pipeline that can effectively and efficiently transforms raw speech into high-quality training data, facilitating the scale-up of the Emilia dataset. The initial version of the Emilia dataset includes over 101k hours of speech data in six languages, featuring a wide variety of in-the-wild real-world speech. Both objective and subjective evaluations confirm the effectiveness of the Emilia dataset. This work aims to advance speech generation towards producing high-quality, spontaneous, and human-like speech, and we invite the research community to utilize Emilia-Pipe for large-scale speech generation.

### v3

@@ 4.3 Multilingual TTS
En AR 6.2% 0.614 14.82
NAR 5.8% 0.581 14.47
Zh AR 4.1% 0.564 35.59
NAR 4.7% 0.560 48.95
De AR 6.8% 0.680 32.72
NAR 13.3% 0.633 44.68
Fr AR 8.2% 0.589 42.92
NAR 17.5% 0.550 48.82
Ja AR 3.6% 0.625 49.42
NAR 10.8% 0.562 49.39
Ko AR 10.9% 0.681 47.93
NAR 15.2% 0.608 63.00
Next, we conduct experiments to evaluate the performance of the multilingual TTS capability of the baselines trained on the full Emilia datasets, encompassing six languages. The test set for En is Emilia-Test. The test set for Zh is from Aishell-3. The test sets for De, Fr, Ja, and Ko are from Common Voice. Each test set contains at least 500 reference samples. Due to space limitations, we only present the objective evaluation results in Table 5, omitting the subjective evaluation results. The findings confirm that both models exhibit strong zero-shot multilingual TTS performance, underscoring the multilingual effectiveness of the Emilia dataset.
@@ 5 Conclusions
In conclusion, this paper introduces Emilia, an extensive, multilingual, and diverse dataset for speech generation, along with Emilia-Pipe, an open-source preprocessing pipeline that can effectively and efficiently transforms raw speech into high-quality training data. The initial version of the Emilia dataset includes over 101k hours of speech data in six languages, featuring a wide variety of in-the-wild real-world speech. Both objective and subjective evaluations demonstrate that training on this diverse dataset can significantly advance speech generation, enabling models to produce more natural, spontaneous, and human-like speech. We have open-sourced both the Emilia dataset and the Emilia-Pipe pipeline, inviting the research community to collaborate on large-scale, spontaneous speech generation initiatives and further drive innovation in this field.